OpenStudy (sally6sa):

Describe how the graph of y=-4f(2(x+3))+6 can be obtained from the graph y=f(x)?

OpenStudy (irishboy123):

for \(y=-4f(2(x+3))+6\) .... via a few **linear* transformations \(\bar y = \dfrac{y - 6}{-4} \) \(\bar x = 2(x+3)\) that kinda stuff , i think, should be good :)

OpenStudy (irishboy123):

so \(\bar y =f (\bar x)\) !!!!